A cubical block of wood weighing 200g has a lead piece fastened underneath. Find the mass of the lead piece which will just allow the block to float in water. Specific gravity of wood is 0.8 and that of lead is 11.3.

Grade:9

2 Answers

Akash Kumar Dutta
11 years ago

Dear Vrushabh,
Let the mass be m,
then to just float Total downward force=bouyant force
also volume of wooden block=m/d=200/.8=250 cm^3
Hence (200 +m)g = (250+m/11.3 ).1.g
Solving we get...54.85 gm (ANS)
Regards.

Chetan Mandayam Nayakar
11 years ago

(m/11.3)+(200/0.8)=200+m